# Exploring Isospin Symmetry Breaking in Exotic Nuclei: High-Precision Mass Measurement of 23Si and Shell-Model Calculations of $T=5/2$ Nuclei

### v2

@@ Abstract
We present a high-precision mass measurement of the proton-rich nucleus 23Si, performed with the LEBIT Penning trap at the Facility for Rare Isotope Beams (FRIB) utilizing the time-of-flight ion cyclotron resonance (TOF-ICR) technique. We determined a mass excess of 23362.9(5.8) keV, which agrees with a recent storage-ring measurement from CSRe but has a factor 20 improved precision. 23Si is hence the nucleus with the most precisely known mass of all nuclei with an isospin projection of $T_{z}=-5/2$. We performed shell-model calculations with the USDC and USDCm Hamiltonians to study binding energy differences and Thomas-Ehrmann shifts in mirror systems with an isospin up to $T=5/2$. Our experimental result and other recently reported masses of neutron-deficient sd-shell nuclei agree well with the theoretical predictions, demonstrating that isospin symmetry breaking in sd-shell nuclei — even at high isospin values — is well described by modern shell-model calculations.
@@
where $a$, $b$ and $c$ are fitting parameters and $\alpha$ describes the spin-parity of the state. The $12T+$ 1 nuclear states of a given isobaric multiplet are also often referred to as isobaric analog states. They have the same isospin $T$, the same mass number $A$ and the same spin-parity $J^{\pi}$ but a different isospin projection $T_{z}$. The IMME is used to predict masses for modeling stellar evolution, as well as to estimate the location of the proton dripline, which is essential to understand the limits of nuclear stability [4]. Moreover, isospin symmetry also plays a significant role in particle physics, especially in testing the unitarity of the Cabibbo-Kobayashi-Maskawa (CKM) matrix [5, 6, 7, 8]. In this way, the study of isospin symmetry breaking is not only relevant for nuclear physics but also deepens our understanding of particle physics and the intricate relationships between the forces that govern both the atomic nucleus and subatomic particles.
In recent years, advances in experimental techniques have provided new data on exotic nuclei far from stability with high isospin values, see Fig. 1. In particular, the masses of several $T_{z}=-5/2$ nuclei have only very recently been published, such as 35Ca [9] in 2023 and 21Al [10], 23Si [11], 27[11] and 31Ar [11] in 2024. These new masses, along with the high-precision mass measurement of 23Si presented in this work, provide a rich data set to explore isospin symmetry breaking in nuclei with high isospin. The mass of 23Si was measured at the Low Energy Beam and Ion Trap (LEBIT) at the Facility for Rare Isotope Beams (FRIB) and improves the previous precision by a factor 20. By comparing the experimental binding energy differences of mirror nuclei up to $T=5/2$ with shell-model calculations [12], we can refine our understanding of the Coulomb force’s role and other isospin symmetry breaking effects and contribute to the broader effort to decode the intricate dynamics of atomic nuclei.
@@
The ions of interest were produced at FRIB by projectile fragmentation. After acceleration to an energy of 290 MeV/u in FRIB’s superconducting Linear Accelerator [16] the primary 28Si28+ beam impinged upon a 17.67 mm thick 12C target creating a wide variety of different ion species including 23Si23+. The ions of interest were separated by the Advanced Rare Isotope Separator (ARIS) [17] from most of the other reaction products, as well as the primary beam, and sent to the Advanced Cryogenic Gas Stopper (ACGS) [18]. The momentum of the purified beam was compressed using a 1004 $\mu$m thick Al wedge at an opening angle of 5 mrad as well as a 7899 $\mu$m thick Al degrader at an angle of 41 degrees in the last energy-dispersive beamline section leading to ACGS. (Note that these settings were optimized for another experiment, in which fully-stripped 23Si was present as an isotone.) Afterwards the beam was stopped in the gas stopper via collisions with helium buffer gas. To guide the ions to the extraction orifice, radio-frequency (RF) carpet surfing [19] was used. After extraction, the ions were passed through a radiofrequency quadrupole acting as a beam cooler and differential pumping barrier, accelerated to a beam energy of 30 keV and mass-separated by a dipole magnet with a mass resolving power of $\approx 1500$. A continuous beam consisting only of 23Si+, 23Al+, 23Mg+ and 23Na+ was then sent towards LEBIT. At LEBIT, the ions were injected into a linear Paul-trap cooler-buncher [20] which contained helium buffer gas to facilitate accumulation, cooling and bunching of the ions. After a cooling time of $\approx 15$ ms, the ions were extracted from the Paul trap as well-defined ion bunches with significantly reduced longitudinal and transversal emittance. Subsequently, they were guided into LEBIT’s 9.4 T hyperbolic Penning trap for further manipulation and analysis [21]. In the Penning trap, the ions were confined in three dimensions by a superposition of a magnetic field $B$ and an electrostatic quadrupole field. The motion of the ion is described by three distinct eigenfrequencies: an axial frequency $\nu_{z}$ and two radial frequencies, $\nu_{-}$ (magnetron frequency) and $\nu_{+}$ (reduced cyclotron frequency). In ideal conditions, the cyclotron frequency $\nu_{c}$ is given by the sum of the two radial frequencies, $\nu_{c}=\nu_{+}+\nu_{-}$ [22]. It is related to the ion’s mass-over-charge ratio $m/q$ via

### v3

@@ Abstract
We present a high-precision mass measurement of the proton-rich nucleus 23Si, performed with the LEBIT Penning trap at the Facility for Rare Isotope Beams (FRIB) utilizing the time-of-flight ion cyclotron resonance (TOF-ICR) technique. We determined a mass excess of 23362.9(5.8) keV, which agrees with a recent storage-ring measurement from CSRe but has a factor of 20 improved precision. 23Si is hence the nucleus with the most precisely known mass of all nuclei with an isospin projection of $T_{z}=-5/2$. We performed shell-model calculations with the USDC and USDCm Hamiltonians to study binding energy differences and Thomas-Ehrmann shifts in mirror systems with an isospin up to $T=5/2$. Our experimental result and other recently reported masses of neutron-deficient sd-shell nuclei agree well with the theoretical predictions, demonstrating that isospin symmetry breaking in sd-shell nuclei — even at high isospin values — is well described by modern shell-model calculations.
@@
where $a$, $b$ and $c$ are fitting parameters and $\alpha$ describes the spin-parity of the state. The $12T+$ 1 nuclear states of a given isobaric multiplet are also often referred to as isobaric analog states. They have the same isospin $T$, the same mass number $A$ and the same spin-parity $J^{\pi}$ but a different isospin projection $T_{z}$. The IMME is used to predict masses for modeling stellar evolution [5, 6, 7], as well as to estimate the location of the proton dripline, which is essential to understand the limits of nuclear stability [4]. Moreover, isospin symmetry and its breaking also play a significant role in particle physics, especially in testing the unitarity of the Cabibbo-Kobayashi-Maskawa (CKM) matrix [8, 9, 10, 11]. In this way, the study of isospin symmetry breaking is not only relevant for nuclear physics but also deepens our understanding of particle physics and the intricate relationships between the forces that govern both the atomic nucleus and subatomic particles.
In recent years, advances in experimental techniques have provided new data on exotic nuclei far from stability with high isospin values, see Fig. 1. In particular, the masses of several $T_{z}=-5/2$ nuclei have only very recently been published, such as 35Ca [12] in 2023 and 21Al [13], 23Si [14], 27[14] and 31Ar [14] in 2024. These new masses, along with the high-precision mass measurement of 23Si presented in this work, provide a rich data set to explore isospin symmetry breaking in nuclei with high isospin. The mass of 23Si was measured at the Low Energy Beam and Ion Trap (LEBIT) at the Facility for Rare Isotope Beams (FRIB) and improves the previous precision by a factor of 20. By comparing the experimental binding energy differences of mirror nuclei up to $T=5/2$ with shell-model calculations [15], we can refine our understanding of the Coulomb force’s role and other isospin symmetry breaking effects and contribute to the broader effort to decode the intricate dynamics of atomic nuclei.
@@
The ions of interest were produced at FRIB by projectile fragmentation. After acceleration to an energy of 290 MeV/u in FRIB’s superconducting Linear Accelerator [19] the primary 28Si14+ beam impinged upon a 17.67 mm thick 12C target creating a wide variety of different ion species including 23Si14+. The ions of interest were separated by the Advanced Rare Isotope Separator (ARIS) [20] from most of the other reaction products, as well as the primary beam, and sent to the Advanced Cryogenic Gas Stopper (ACGS) [21]. The momentum of the purified beam was compressed using a 1004 $\mu$m thick Al wedge at an opening angle of 5 mrad as well as a 7899 $\mu$m thick Al degrader at an angle of 41 degrees in the last energy-dispersive beamline section leading to ACGS. These settings were optimized for another experiment, in which fully-stripped 23Si was present as an isotone. Afterwards the beam was stopped in the gas stopper via collisions with helium buffer gas. To guide the ions to the extraction orifice, radio-frequency (RF) carpet surfing [22] was used. After extraction, the ions were passed through a radiofrequency quadrupole acting as a beam cooler and differential pumping barrier, accelerated to a beam energy of 30 keV and mass-separated by a dipole magnet with a mass resolving power of $\approx 1500$. A continuous beam consisting only of 23Si+, 23Al+, 23Mg+ and 23Na+ was then sent towards LEBIT. At LEBIT, the ions were injected into a linear Paul-trap cooler-buncher [23] which contained helium buffer gas to facilitate accumulation, cooling and bunching of the ions. After a cooling time of $\approx 15$ ms, the ions were extracted from the Paul trap as well-defined ion bunches with significantly reduced longitudinal and transversal emittance. Subsequently, they were guided into LEBIT’s 9.4 T hyperbolic Penning trap for further manipulation and analysis [24]. In the Penning trap, the ions were confined in three dimensions by a superposition of a magnetic field $B$ and an electrostatic quadrupole field. The motion of the ion is described by three distinct eigenfrequencies: an axial frequency $\nu_{z}$ and two radial frequencies, $\nu_{-}$ (magnetron frequency) and $\nu_{+}$ (reduced cyclotron frequency). In ideal conditions, the cyclotron frequency $\nu_{c}$ is given by the sum of the two radial frequencies, $\nu_{c}=\nu_{+}+\nu_{-}$ [25]. It is related to the ion’s mass-over-charge ratio $m/q$ via